OpenStudy (anonymous):

solve the equation using the zero product property (2x+6)(3x-6)=0 a. 3 and 2 b. -3 and -2 c. -3 and 2 d. 3 and -2

OpenStudy (butterflydreamer):

If we use the zero product property it basically means this: We have (2x + 6) (3x -6) = 0 Meaning that 2x + 6 = 0 AND 3x - 6 = 0 So now for each case solve for x :)

OpenStudy (anonymous):

i'm not sure how to solve for x.

OpenStudy (butterflydreamer):

oh no problem :) We'll work it out step by step: